OpenStudy (anonymous):

suppose you wanted to find two consecutive positive even integers whose product is 440. How would you go about finding these two integers?

OpenStudy (anonymous):

so if we let one of this numbers be the variable x can you tell me what the second number is in therms of x?

OpenStudy (anonymous):

No because I don't know how to do this.

OpenStudy (anonymous):

okay lets break this down even further if just stick with me here. What is the next even integer after 2?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now whats the next even integer after 40?

OpenStudy (anonymous):

42??

OpenStudy (anonymous):

or 60?

OpenStudy (anonymous):

...42 I don't know where you got 60 from. Remember a even number is a number divisible by two so if I'm at an even number how much should I add to get to the next one?

OpenStudy (anonymous):

So 42 is the first integer for this problem?

OpenStudy (anonymous):

no no no I\m just trying to help you get the first step in solving this on your own

OpenStudy (anonymous):

lets put it another way.... whats 4-2=? whats 42-40=? whats 112-110=?

OpenStudy (anonymous):

I'm so frustrated...

OpenStudy (anonymous):

its okay I'll work through it with you...sometimes these math problems take time to see the first step. I won't just give you the step cause helping you through it will help you with other problems further on down the course.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

so the problems I gave you all involve a number and their consecutive even integer and should help you see the pattern they have. Try to solve them.

OpenStudy (anonymous):

killer....please do not give him the answer

OpenStudy (anonymous):

to solve this all you have to do is set the equation with a variable the equation would be x+(x+2)=440

OpenStudy (anonymous):

...in other words what I was trying to get him to figure out on his own,,,,

OpenStudy (anonymous):

oh im sorry about that

OpenStudy (anonymous):

yeah well you can help him now

OpenStudy (anonymous):

I am a HER and thank you killerdime20

OpenStudy (anonymous):

your welcome and do you still need help?